Helloooo~ Can you help me with these problem? :)
How many ounces of pure silver should be added to 180 ounces, 75% pure, to obtain a mixture of 80% pure?
Thaaaank yoou in advance. Good day! :)

Answer by ptaylor(2198)

Let x=amt of pure silver needed
Now we know that the amount of PURE silver that exists before the mixture takes place has to equal the amount of PURE silver that exists after the mixture takes place:
Pure silver before the mixture takes place: x+180*0.75=x+135
Pure silver after the mixture takes place: (180+x)*0.80=144+0.80x
Sooooooooooooo:
x+135=0.80x+144 subtract 0.80x and also 135 from both sides
x-0.80x+135-135=0.80x-0.80x+144-135 collect like terms
0.20x=9
x=45 oz-----------------amt of pure silver needed
CK
45+135=144+0.80*45
180=180
Hope this helps---ptaylor